OpenStudy (anonymous):

Find the GCF of the list of monomials. 64x^8y^3, 40x^6y^8

OpenStudy (campbell_st):

looks like 8x^6y^3

OpenStudy (anonymous):

hod did you get 8x^6?

OpenStudy (anonymous):

how*

OpenStudy (campbell_st):

largest common factor of 64 and 40 is 8.... largest common factor of x^8 and x^6 is x^6 laegest common factor of y^3 and y^8 is y^3 mutliply them together

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

64x^8y^3, 40x^6y^8 64 = 8 * 8 40 = 8 * 5 -> LCM = 8 x ^ 8 = x^6 * x^2 -> LCM: x^2 y^8 = y^3 * y^5 --> LCM : y^3